OpenStudy (anonymous):

Line h passes through (8, 5) and (12, 11). Graph the line perpendicular to h that passes through the point (7, 10). What is the slope of the perpendicular line?

OpenStudy (shamim):

the slope of ur line h is\[m _{1}=\frac{ 11-5 }{ 12-8 }=\frac{ 6 }{ 4 }=\frac{ 3 }{ 2 }\]

OpenStudy (shamim):

if the slope of perpendicular line is\[m _{2}\]

OpenStudy (shamim):

then\[m _{1}m _{2}=-1\]

OpenStudy (shamim):

\[\frac{ 3 }{ 2 }m _{2}=-1\]

OpenStudy (shamim):

\[m _{1}=-\frac{ 2 }{ 3 }\]

OpenStudy (shamim):

so the equation of ur perpendicular line is \[y-10=-\frac{ 2 }{ 3 }(x-7)\]

OpenStudy (shamim):

now solve it

OpenStudy (anonymous):

so the slope of the perpendicular line is -2/3?

OpenStudy (shamim):

did u get it

OpenStudy (shamim):

ya

OpenStudy (anonymous):

ohh thanks ur a lifesaver

OpenStudy (shamim):

welcome

OpenStudy (shamim):

u can give me a medal

OpenStudy (anonymous):

from where ?

OpenStudy (anonymous):

lsn

OpenStudy (shamim):

i got medal

OpenStudy (anonymous):

do u do h.ws?

OpenStudy (anonymous):

we can do some business if u have time

OpenStudy (shamim):

anyway which grade u read?

OpenStudy (anonymous):

i would get u a job with some friends

OpenStudy (shamim):

of course i hv enough time

OpenStudy (anonymous):

than give me ur email

OpenStudy (anonymous):

my name is adel i will send u an email from my email

OpenStudy (shamim):

md_habib33@yahoo.com

OpenStudy (shamim):

ok u can send email

OpenStudy (anonymous):

where do u study ? and what degree?

OpenStudy (shamim):

we can communicate by facebook too

OpenStudy (shamim):

i m msc of physics

OpenStudy (shamim):

assistant professor of physics

OpenStudy (shamim):

can i get ur facebook id

OpenStudy (anonymous):

ya

OpenStudy (shamim):

i wanna add u in facebook

OpenStudy (anonymous):

Adel Balubaid

OpenStudy (shamim):

ok

OpenStudy (anonymous):

or search by email : adel.balubaid2323@gmail.com

OpenStudy (anonymous):

did u find me ?

OpenStudy (shamim):

ya

OpenStudy (shamim):

sent add request and message

OpenStudy (shamim):

u can accept my request and we can chat in facebook abt ur business

OpenStudy (anonymous):

i didnt get ur msg

OpenStudy (anonymous):

ur name mohammed shamim ?

OpenStudy (shamim):

ya

OpenStudy (anonymous):

i added u

OpenStudy (shamim):

u can search by my email address too in facebook

OpenStudy (anonymous):

yeah i made the request

OpenStudy (anonymous):

never mind i will email u

OpenStudy (shamim):

ok

OpenStudy (shamim):

i accept ur friend request